# Hybrid Many-Objective Optimization in Probabilistic Mission Design for Compliant and Effective UAV Routing

### v1

@@ 5.1. Summary
We have presented routing for Unmanned Aerial Vehicles (UAV) through Hybrid Multi-Objective Optimization within Probabilistic Mission Design (ProMis). As such, the illustrated methods enable encoding physical objectives with airspace regulations. While the former is modeled as cost functions over UAV paths or points in navigation space, the latter is facilitated through probabilistic first-order logic and spatial reasoning parameterized by Statistical Relational Maps (StaR Map).
As an integral part of the framework, we have presented the Constitutional Language (CoLa) as a unified description language to formalize an agent’s navigational objectives and constraints in a single model. To this end, CoLa lets the user define high-level mission parameters such as the aviation license of the employed operator or the time of day at which the mission is conducted. These parameters hereby span up a space for Explanation and Optimization steps beyond path optimization, i.e., showing which part of the setting leads to a denial of Clearance for the intended path or what setting would yield ideal circumstances from a legal viewpoint. Furthermore, CoLa allows for encoding aviation rules across a StaR Map parameterized space as scalar fields of identically and independently distributed probabilities of satisfied constraints under uncertainty.
We have shown how Hybrid Multi-Objective Optimization in ProMis enables large-scale, physically, and legally informed UAV routing in a statistical evaluation of the framework’s application in Paris, France. Based on low-cost, crowd-sourced map data provided by OpenStreetMap, a set of physical objectives about the agent’s path and states, as well as a simplified model of legal and safety constraints, our experiments demonstrate how these methods can create networks of compliant paths over vast human-inhabited spaces. Along these lines, we have illustrated how this setting allows for a meta-analysis of the satisfiable airspace, i.e., how to decide Clearance for individual paths, Explain the impact of mission parameters, and Optimize for the most suitable ones with regard to compliance.
@@ Acknowledgments
The Eindhoven University of Technology authors received support from their Department of Mathematics and Computer Science and the Eindhoven Artificial Intelligence Systems Institute. Map data © OpenStreetMap contributors, licensed under the Open Database License (ODbL) and available from https://www.openstreetmap.org. Map styles © Mapbox, licensed under the Creative Commons Attribution 3.0 License (CC BY 3.0) and available from https://github.com/mapbox/mapbox-gl-styles. Cell tower data © OpenCellID, licensed under the Creative Commons Attribution ShareAlike 4.0 License (CC BY-SA 4.0) and available under https://wiki.opencellid.org.

### v2

@@ 5.1. Summary
We have presented a compliant and effective routing for UAVs through Hybrid Multi-Objective Optimization within Probabilistic Mission Design (ProMis). As such, our framework enables UAV operators to jointly encode physical objectives for effectiveness and airspace regulations for compliance. While the former is modeled as cost functions over UAV paths or points in navigation space, the latter is facilitated through probabilistic first-order logic and spatial reasoning parameterized by Statistical Relational Maps (StaR Map).
As an integral part of our framework, we have presented the Constitutional Language (CoLa) as a unified description language to formalize an agent’s navigational objectives and constraints in a single model. To this end, CoLa lets the user define high-level mission parameters such as the aviation license of the employed operator or the time of day at which the mission is conducted. These parameters hereby span up a space for Explanation and Optimization steps beyond path optimization, i.e., showing which part of the setting leads to a denial of Clearance for the intended path or what setting would yield ideal circumstances from a legal viewpoint. Furthermore, CoLa allows for encoding aviation rules across a StaR Map parameterized space as scalar fields of identically and independently distributed probabilities of satisfied constraints under uncertainty.
We have shown how Hybrid Multi-Objective Optimization in ProMis enables large-scale, physically and legally informed UAV routing in a statistical evaluation of the framework’s application in Paris, France. Based on low-cost, crowd-sourced map data provided by OpenStreetMap, a set of physical objectives about the agent’s path and states, as well as a simplified model of legal and safety constraints, our experiments demonstrate how these methods can create networks of compliant paths over vast human-inhabited spaces. Along these lines, we have illustrated how this setting allows for a meta-analysis of the satisfiable airspace, i.e., how to decide Clearance for individual paths, Explain the impact of mission parameters, and Optimize for the most suitable ones with regard to compliance. In our experiments, this reduces the path rejection area over curve, e.g., by approximately in the case of knee point paths.
@@ Acknowledgments
Simon Kohaut and Nikolas Hohmann gratefully acknowledge the financial support from Honda Research Institute Europe (HRI-EU). The Eindhoven University of Technology authors received support from their Department of Mathematics and Computer Science and the Eindhoven Artificial Intelligence Systems Institute. Map data © OpenStreetMap contributors, licensed under the Open Database License (ODbL) and available from https://www.openstreetmap.org. Map styles © Mapbox, licensed under the Creative Commons Attribution 3.0 License (CC BY 3.0) and available from https://github.com/mapbox/mapbox-gl-styles. Cell tower data © OpenCellID, licensed under the Creative Commons Attribution ShareAlike 4.0 License (CC BY-SA 4.0) and available under https://wiki.opencellid.org.